OpenStudy (anonymous):

find all solutions of the equation tan^5x-9tan x=0 the answer is Akpi where k is any integer the constant A is ?

OpenStudy (athe):

try \[\tan^5{x}-9\tan{x}=0 \rightarrow \tan{x}\cdot(\tan^4{x}-9)=0\] eqv: \[\tan{x}=0\]or\[\tan{x}=\pm\sqrt{3}\]

OpenStudy (athe):

and solve this elementary equations

OpenStudy (anonymous):

i did and the answer is 0.3333 but im getting something different

OpenStudy (athe):

\[\tan{x}=0 \rightarrow x=\pi\cdot n, n \in Z\]

OpenStudy (athe):

\[\tan{x}=\pm\sqrt{3} \rightarrow x=\pm \frac{ \pi }{ 3 } +\pi\cdot n, n \in Z\]

OpenStudy (anonymous):

thank you

OpenStudy (athe):

ok